Solve each system by substitution. Check your solution.
19.
a-1.2b=-3
0.2b+0.6a=12

Answer by checkley77(12844)

a-1.2b=-3 or a=1.2b-3
0.2b+0.6a=12
.2b+.6(1.2b-3)=12
.2b+.72b-1.8=12
.92b=12+1.8
.92b=13.8
b=13.8/.92
b=15 ans.
a=1.2*15-3
a=18-3
a=15 ans.
Proof:
.2*15+.6*15=12
3+9=12
12=12